Internet Engineering Task Force

INTERNET-DRAFT  Schulzrinne
draft-schulzrinne-tftp-url-01.ps  Columbia U.

November 19, 2001 Expires: June 2002

The tftp URL Scheme

Status of this Memo

This document is an Internet-Draft and is in full conformance with all provisions of Section 10 of RFC2026.

Internet-Drafts are working documents of the Internet Engineering Task Force (IETF), its areas, and its working groups. Note that other groups may also distribute working documents as Internet-Drafts.

Internet-Drafts are draft documents valid for a maximum of six months and may be updated, replaced, or obsoleted by other documents at any time. It is inappropriate to use Internet-Drafts as reference material or to cite them other than as “work in progress.”

The list of current Internet-Drafts can be accessed at http://www.ietf.org/ietf/1id-abstracts.txt To view the list Internet-Draft Shadow Directories, see http://www.ietf.org/shadow.html.

Copyright Notice

Copyright (c) The Internet Society (2001). All Rights Reserved.

Abstract

This document defines the tftp URL according to the rules in RFC 2717.

1 URL Scheme Name

The scheme name is “tftp”.

2 URL Scheme Syntax

The TFTP URL scheme is used to designate files and directories on Internet hosts accessible using the TFTP protocol [1].

TFTP URLs follow the common Internet scheme syntax described in Section 3.1 of RFC 1738 [2]. If “:<port>” is omitted, the port defaults to 69. The “user” and “password” components must not be used. No relative forms are supported.

An example of a TFTP URL is

  tftp://tftp.example.com/directory/file.txt

3 Character Encoding Considerations

RFC 1350 only allows file names consisting of “netascii” characters [3]. The host part of the URL has the same encoding considerations as described in RFC 1738 [2].

INTERNET-DRAFT  draft-schulzrinne-tftp-url-01.ps  November 19, 2001

4 Intended Usage

The file identified is retrieved (read) from the host and file name specified in the URL, using TFTP read requests (RRQ).

5 Protocols which Use This URL Scheme Name

TFTP (STD 33, RFC 1350 [1])

6 Interoperability considerations

TFTP mail mode and Netascii modes are not supported. (RFC 1350 states that “mail mode is obsolete and should not be implemented or used.”)

7 Security Considerations

The security considerations of RFC 1350 [1] apply. This scheme is suitable only for files that are considered public, similar to anonymous ftp or public HTTP objects. [TBD: Explain any differences to anon. ftp.] For security reasons, this URL does not support the TFTP WRQ (write) mode.

8 Author Address

Henning Schulzrinne Dept. of Computer Science Columbia University 1214 Amsterdam Avenue New York, NY 10027 USA electronic mail: schulzrinne@cs.columbia.edu

References

[1] K. Sollins, “The TFTP protocol (revision 2),” Request for Comments 1350, Internet Engineering Task Force, July 1992.

[2] T. Berners-Lee, L. Masinter, and M. McCahill, “Uniform resource locators (URL),” Request for Comments 1738, Internet Engineering Task Force, Dec. 1994.

[3] J. Postel, “Telnet protocol specification,” Request for Comments 764, Internet Engineering Task Force, June 1980.

Full Copyright Statement

Copyright (c) The Internet Society (2001). All Rights Reserved.

This document and translations of it may be copied and furnished to others, and derivative works that comment on or otherwise explain it or assist in its implementation may be prepared, copied, published and distributed, in whole or in part, without restriction of any kind, provided that the above copyright notice and this paragraph are included on all such copies and derivative works. However, this document itself may not be modified in any way, such as by removing the copyright notice or references to the Internet Society or other Internet organizations, except as needed for the purpose of developing Internet standards in which case the procedures for copyrights defined in the Internet Standards process must be followed, or as required to translate it into languages other than English.

The limited permissions granted above are perpetual and will not be revoked by the Internet Society or its successors or assigns.

This document and the information contained herein is provided on an ”AS IS” basis and THE IN-TERNET SOCIETY AND THE INTERNET ENGINEERING TASK FORCE DISCLAIMS ALL WAR-RANTIES, EXPRESS OR IMPLIED, INCLUDING BUT NOT LIMITED TO ANY WARRANTY THAT THE USE OF THE INFORMATION HEREIN WILL NOT INFRINGE ANY RIGHTS OR ANY IMPLIED WARRANTIES OF MERCHANTABILITY OR FITNESS FOR A PARTICULAR PURPOSE.